# Dejavu: Towards Experience Feedback Learning for Embodied Intelligence

### v2

@@ 3.4 Deployment Retrieval and Experience Growth
Goal and key differences. At deployment, EFN recalls and reuses prior experiences while not updating the base policy’s or EFN’s weights. The inference pipeline mirrors training but differs in three places. First, retrieval is task-filtered: we restrict matches to rollouts whose instruction embeddings are close to the current task. Second, we prioritize efficient rollouts: shorter trajectories receive higher selection priority because they typically contain fewer redundant actions and lead to faster completion. Third, we grow the bank online: after a rollout finishes, its steps are inserted into the bank so that future episodes can recall them. Thus, post-deployment adaptation comes purely from memory growth, not gradient updates.
Instruction-filtered candidate set. Given a task description, we compute an instruction embedding with the VLA’s language encoder, denoted $\ell^{\star}$. We compare $\ell^{\star}$ to all stored rollout-level embeddings $\{\ell_{\tau_{j}}\}$ with cosine similarity and select the top-$n$ rollouts:
@@
All step-level entries from these rollouts form the candidate experience set $\mathcal{C}$, which is the only pool we retrieve from during this episode.
Step-wise retrieval with efficiency prior. At step $t$, we form a visual query $\mathbf{q}_{t}=\mathbf{u}(\mathbf{F}_{t})$ using the mean–max fusion $\mathbf{u}(\cdot)$ defined previously. Each candidate $i\in\mathcal{C}$ has a key $\mathbf{k}_{i}$, a successor feature $\hat{\mathbf{F}}^{+}_{i}$, and belongs to a rollout $\rho(i)$ with total length $L_{\rho(i)}$. We score candidates by combining semantic similarity and an efficiency prior that favors shorter rollouts. Let $s_{i}=\cos(\mathbf{q}_{t},\mathbf{k}_{i})$ and define a normalized length prior $g(L_{\rho(i)})\in[0,1]$ that decreases with $L_{\rho(i)}$ (e.g., $g(L)=\exp[-\beta\,L/\bar{L}]$ with temperature $\beta$ and reference length $\bar{L}$). The combined score is
@@
Online experience growth. After each episode, we insert the resulting rollout into the experience bank: the episode-level instruction embedding $\ell^{\star}$ is stored together with all step tuples $\big(\mathbf{F}_{t},\mathbf{k}_{t},\mathbf{a}_{t}^{(0)}\big)$ computed under the same mean–max keying scheme. At deployment time, however, we only append rollouts that successfully reach the goal, unlike the training phase which admits both successful and failed trajectories (see Appendix D for discussion). These successful episodes serve as high-quality references for future retrieval. The initial offline bank may still contain near-successes or even failures, as also noted in Appendix D. When operating under an experience budget, standard retention strategies (e.g., reservoir sampling or recency-aware replacement) can be applied without modifying the retrieval or learning mechanisms described above.
Table 1: Performance on Libero Dataset reported by the average results evaluated across three seeds.
@@
Method Spatial Object Goal Long Average Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Base: OpenVLA OpenVLA [openvla] 84.7 119.5 88.4 163.7 79.2 121.5 53.7 275.9 76.5 160.2 +kNN-RAG 82.3 125.1 85.6 171.2 74.8 128.3 50.1 282.7 73.2 166.4 +ResAct [resact] 86.1 117.3 89.7 160.5 82.5 119.7 58.4 271.5 79.2 158.6 +R2A [r2a] 87.5 114.8 91.1 157.2 84.6 116.9 63.2 265.3 81.6 156.3 +GC-TTT [gcttt] 85.2 121.9 87.9 165.8 80.7 123.6 55.9 277.1 77.4 162.8 +EFN (Vol=300) 88.5 115.4 91.3 158.8 85.7 117.0 72.1 267.2 84.4 160.0 +EFN (Vol=1000) 89.9 109.0 92.2 156.1 89.2 115.2 76.5 261.9 87.0 156.7 Base: UniVLA UniVLA [Bu2025UniVLA] 96.5 112.7 96.8 159.0 95.6 124.9 92.0 264.5 95.2 164.2 +kNN-RAG 94.2 118.3 95.7 165.1 93.1 128.6 87.5 270.3 92.6 168.9 +ResAct [resact] 97.1 109.6 97.4 154.7 96.2 122.1 89.8 260.7 95.1 160.0 +R2A [r2a] 97.5 107.2 97.8 152.3 96.6 120.5 91.3 257.9 95.8 158.0 +GC-TTT [gcttt] 96.8 110.1 97.1 156.8 95.9 123.4 89.1 266.1 94.7 162.1 +EFN (Vol=300) 97.7 103.4 97.9 151.4 97.2 120.1 93.7 253.8 96.6 156.2 +EFN (Vol=1000) 98.2 102.1 98.2 145.8 97.6 117.8 94.6 242.5 97.2 151.3 Base: GO-1 GO-1 [go1] 96.3 111.6 97.4 160.2 95.6 118.9 89.3 269.1 94.7 163.1 +kNN-RAG 94.1 116.8 95.9 167.5 93.4 122.7 85.1 274.9 92.1 168.0 +ResAct [resact] 96.8 108.9 97.6 156.1 96.1 119.8 88.7 264.7 94.8 160.3 +R2A [r2a] 97.3 106.5 97.9 153.6 96.5 118.2 90.2 261.5 95.5 158.1 +GC-TTT [gcttt] 96.6 113.2 97.2 159.5 95.8 121.1 88.9 267.3 94.6 163.3 +EFN (Vol=300) 97.5 107.8 98.1 154.5 96.8 115.6 91.2 258.5 95.9 157.5 +EFN (Vol=1000) 98.1 105.3 98.5 151.7 97.3 114.1 92.8 253.2 96.7 154.8
@@
Figure 6: Visualization of EFN’s residual corrections. Left: current observation; Middle: retrieved experience frame; Right: corrected (translucent) vs. base action. Even when the retrieved scene differs from the current one (top row), EFN still extracts actionable guidance.
@@
Table 2: Real-world performance on the AgiBot-G1 platform with the GO-1 policy [go1].
@@
Method BottlePlace ShelfSort StockLift DrawerStore Average Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ GO-1 [go1] 47.3 407.0 34.0 442.1 16.0 389.2 5.3 728.9 25.8 491.8 +kNN-RAG 35.3 435.2 25.3 468.7 12.0 405.1 0.0 752.3 18.2 515.3 +ResAct [resact] 52.7 418.9 41.3 431.4 22.7 396.8 8.0 715.6 31.2 490.7 +R2A [r2a] 61.3 399.1 51.3 415.7 31.3 372.5 14.0 689.2 39.5 469.1 +GC-TTT [gcttt] 49.3 412.5 36.7 448.9 18.7 393.7 0.0 735.1 26.2 497.6 +EFN (Vol=300) 69.3 382.6 54.7 410.5 42.0 355.1 37.3 668.3 50.8 454.1 +EFN (Vol=1000) 82.0 364.2 74.7 392.9 65.3 339.9 58.7 643.4 70.2 435.1
@@ 4 Experiments
We evaluate EFN on the Libero benchmark [Liberoneurips2023] with three pretrained VLAs: OpenVLA [openvla], UniVLA [univla], and GO-1 [go1]. All models share the same visual preprocessing (center crop, resize to $256$). UniVLA provides $256$ visual tokens and $4$ latent action tokens per step; EFN takes the current tokens and retrieved tokens, predicts a residual in latent space, and adds it to the base latent action, which is decoded by the frozen action head. Training details (optimizer, replay, schedules) are in Appendix B.
Evaluation protocol. At test time the policy is deterministic by applying $\tanh$ to the actor mean. We report success rate and, conditional on success, the average number of steps (lower is better) with horizon H=320H{=}320. The experience bank used at evaluation mirrors training and is queried via cosine nearest neighbors over pooled visual embeddings. After each successful rollout we append the episode to the bank and prioritise shorter successful episodes; no gradient updates are performed during evaluation.
Baselines and variants. We compare each backbone with four baselines and our EFN variants, chosen to instantiate the three paradigms from Sec. 2. kNN-RAG is a non-parametric retrieval-only controller that retrieves the nearest past step in the shared bank and directly copies its action. ResAct [resact] is a residual-only baseline that learns an additive correction on top of the frozen VLA action but does not use retrieval or similarity-shaped rewards. R2A [r2a] adapts retrieval-augmented RL to VLAs: the policy is conditioned on retrieved trajectories and optimized end-to-end, updating the backbone. GC-TTT [gcttt] is a goal-conditioned test-time training baseline that updates the backbone online using retrieved rollouts. A more detailed baseline setting is described in Appendix B.
@@ 4.2 Benchmark and Analysis
Real-world results. Table 2 reports results on AgiBot-G1 with GO-1 on four real-world tasks: BottlePlace, ShelfSort, StockLift, and DrawerStore (from easy to hard), where the detailed settings of these four tasks are provided in Appendix B. The trends mirror simulation: kNN-RAG consistently hurts performance; ResAct and R2A offer moderate gains; GC-TTT is unstable and fails on the hardest task. EFN achieves the highest average success with fewer steps, and the gap between Vol=300 and Vol=1000 is smaller than in simulation, suggesting that a moderate bank already covers the main task variations. All EFN results are obtained with a frozen GO-1 backbone; adaptation arises solely from the experience bank and the residual head, whereas R2A and GC-TTT rely on gradient-based updates. We include visualizations of real-world experiments in Appendix F.
@@ 4.3 Ablation Study
We further evaluate four controlled variants on LIBERO (Figure 7) with raw data provided in Appendix Table 6.
Effect of SAC. Removing SAC (w/o SAC) and using a value-only critic harms both success and efficiency, supporting the choice of entropy-regularised optimisation.
@@
Figure 7: Ablation studies of EFN on LIBERO Environment.
Table 3: Inference-time efficiency comparison on real-world deployment. Change is computed relative to the Base GO-1.
@@
Effect of instruction. Disabling instruction-level filtering (w/o instr) degrades the performance, indicating that instruction-aware retrieval is important.
Effect of anti-idle. Without the anti-idling term (w/o anti-idle) the performance reduces, indicating that idleness penalties make a difference.
Effect of residual-action. As analyzed in Appendix E, replacing the residual correction with direct action prediction degrades performance, which we attribute to discarding the base policy’s generalization ability and overburdening the lightweight actor with modeling the full action space.
Inference time. Interestingly, EFN improves both performance and efficiency. Although each step introduces a modest 4.2% latency overhead, EFN reduces the average episode length by 11.5%, resulting in an overall 7.9% reduction in total episode time. In other words, EFN achieves higher success rates while running more efficiently.
@@ 5 Conclusion
We introduced Experience Feedback Network, a retrieval-conditioned residual module that augments a frozen vision–language–action policy with an experience bank. EFN retrieves task-relevant transitions in a joint vision–language space and uses similarity-shaped reinforcement learning to train a small residual controller, enabling post-deployment adaptation through memory updates rather than backbone finetuning. On LIBERO and on a real-world AgiBot-G1 platform, EFN consistently outperforms retrieval-only, residual-only, retrieval-augmented RL, and test-time training baselines under the same backbone and interaction budget. We add a detailed discussion in Appendix D.
Appendix
@@ Appendix A Overview
Thank you for reading the Appendix for our research. This Appendix is organized as the following sections. Section B describes the detailed settings of our method, compared baselines, and experiments. Section C elaborates on the rationale, motivation, and intuition of designing our EFN. Section D includes a discussion about our method and this field. Section E includes supplementary results by more experiments and analysis. Section F demonstrates the performance of our EFN with more visualizations.
@@ Appendix B Details of Method and Experiment
This Section focuses on the detailed design, and you can refer to Section C behind such design.
@@ B.1 Experience Bank Implementation Details
$E_{m,n}=\big(\ell_{\tau_{m}},\,\mathbf{F}_{m,n},\,\mathbf{k}_{m,n},\,\mathbf{a}^{(0)}_{m,n}\big)$, (17)
where $\ell_{\tau_{m}}$ is the instruction embedding for rollout $\tau_{m}$, $\mathbf{F}_{m,n}\in\mathbb{R}^{L\times C}$ is the VLA vision encoder feature map for frame $s_{m,n}$, $\mathbf{k}_{m,n}\in\mathbb{R}^{d_{k}}$ is the mean–max fused retrieval key defined in Sec. 3.2, and $\mathbf{a}^{(0)}_{m,n}$ is the raw action produced by the frozen base policy at that step.
In our implementation the bank is stored as a line-based JSON file plus accompanying feature files, where each JSON line encodes one step and includes a rollout identifier, a step index, and paths to the visual and action features. The corresponding visual feature file stores the precomputed retrieval key $\mathbf{k}_{m,n}$ as a $d_{k}$-dimensional vector (e.g., visual_embed_meanmax with $d_{k}=4096$), together with any additional visual features that may be needed. The action file stores the base policy’s latent action representation (e.g., a $4\times d_{a}$ tensor of action tokens) and the associated discrete token IDs (generated_ids) from which the continuous control command $\mathbf{a}^{(0)}_{m,n}$ is decoded.
Figure 8: Cosine similarities (left bars) and corresponding softmax sampling probabilities (right bars) over the top candidates in the experience bank. Higher similarity induces higher sampling probability while still preserving exploration.
@@
$\texttt{index}:(\texttt{rollout\_id},\,\texttt{step\_idx})\mapsto E_{m,n},$ (18)
@@
$\mathbf{K}\in\mathbb{R}^{N\times d_{k}},$ (19)
@@
$E_{m,n+1}=\texttt{index}\big(\texttt{rollout\_id}=m,\;\texttt{step\_idx}=n+1\big).$ (20)
@@
Given a current observation $s_{t}$, we compute the vision encoder feature map $\mathbf{F}_{t}$ and derive a query key $\mathbf{q}_{t}$ via the same mean–max fusion and $\ell_{2}$-normalization as used for $\mathbf{k}_{m,n}$ (Sec. 3.2). We then compute cosine similarities between $\mathbf{q}_{t}$ and all stored keys:
$s_{i}=\cos(\mathbf{q}_{t},\mathbf{k}_{i}),\quad i\in\{1,\dots,N\}.$ (21)
we detail the top-$k$ sampling scheme described in the main text: select a shortlist $\mathcal{N}_{k}(\mathbf{q}_{t})$ of the $k$ highest-similarity keys and sample an index using a temperature-scaled softmax,
$p(i\mid\mathbf{q}_{t})=\frac{\exp\big(s_{i}/\tau\big)}{\sum_{j\in\mathcal{N}_{k}(\mathbf{q}_{t})}\exp\big(s_{j}/\tau\big)},\quad i\in\mathcal{N}_{k}(\mathbf{q}_{t}),$ (22)
with a fixed temperature $\tau>0$ (we use $\tau=0.05$ in all experiments). As shown in Figure 8, this “retrieve-then-sample” design encourages exploration among near-matches while still biasing toward the most semantically similar experiences.
@@
During offline pretraining of EFN, the experience bank is constructed from a mixture of successful, near-successful, and occasionally failed rollouts, as discussed in Appendix D. At each environment step, EFN queries the bank, obtains a reference step $E_{m,n}$ and its successor $E_{m,n+1}$, and uses the successor observation as the semantic target for the similarity-shaped progress reward. At deployment, the bank continues to grow: we append only the trajectories that reach the goal, thereby gradually refining the set of available references without ever updating the frozen VLA backbone weights.

### v3

@@ 3.4 Deployment Retrieval and Experience Growth
Goal and key differences. At deployment, EFN recalls and reuses prior experiences while not updating either the base policy weights or its own. The inference pipeline mirrors training but differs in three ways. First, retrieval is task-filtered: we restrict matches to rollouts whose instruction embeddings are close to the current task. Second, we prioritize efficient rollouts: shorter trajectories receive higher selection priority because they typically contain fewer redundant actions and lead to faster completion. Third, we grow the bank online: after a rollout finishes, its steps are inserted into the bank so that future episodes can recall them. Post-deployment adaptation therefore comes purely from memory growth, not gradient updates.
Instruction-filtered candidate set. Given a task description, we compute an instruction embedding with the VLA’s language encoder, denoted $\ell^{\star}$. We compare $\ell^{\star}$ with all stored rollout-level embeddings $\{\ell_{\tau_{j}}\}$ using cosine similarity and select the top-$n$ rollouts:
@@
All step-level entries from these rollouts form the candidate experience set $\mathcal{C}$, which is the only pool from which we retrieve during this episode.
Step-wise retrieval with efficiency prior. At step $t$, we form a visual query $\mathbf{q}_{t}=\mathbf{u}(\mathbf{F}_{t})$ using the mean–max fusion $\mathbf{u}(\cdot)$ defined previously. Each candidate $i\in\mathcal{C}$ has a key $\mathbf{k}_{i}$, a successor feature $\hat{\mathbf{F}}^{+}_{i}$, and belongs to a rollout $\rho(i)$ with total length $L_{\rho(i)}$. We score candidates by combining semantic similarity with an efficiency prior that favors shorter rollouts. Let $s_{i}=\cos(\mathbf{q}_{t},\mathbf{k}_{i})$ and define a normalized length prior $g(L_{\rho(i)})\in[0,1]$ that decreases with $L_{\rho(i)}$ (e.g., $g(L)=\exp[-\beta\,L/\bar{L}]$ with temperature $\beta$ and reference length $\bar{L}$). The combined score is
@@
Online experience growth. After each episode, we insert the resulting rollout into the experience bank: the episode-level instruction embedding $\ell^{\star}$ is stored together with all step tuples $\big(\mathbf{F}_{t},\mathbf{k}_{t},\mathbf{a}_{t}^{(0)}\big)$ computed under the same mean–max keying scheme. At deployment time, however, we append only rollouts that successfully reach the goal, unlike the training phase, which may also admit near-successful or failed trajectories (Appendix SectionD.1). These successful episodes serve as high-quality references for future retrieval. When operating under an experience budget, standard retention strategies, such as reservoir sampling or recency-aware replacement, can be applied without modifying the retrieval or learning mechanisms described above.
Table 1: Performance on the LIBERO benchmark, averaged over three seeds.
@@
Method Spatial Object Goal Long Average Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Base: OpenVLA OpenVLA [18] 84.7 119.5 88.4 163.7 79.2 121.5 53.7 275.9 76.5 160.2 +kNN-RAG 82.3 125.1 85.6 171.2 74.8 128.3 50.1 282.7 73.2 166.4 +ResAct [29] 86.1 117.3 89.7 160.5 82.5 119.7 58.4 271.5 79.2 158.6 +R2A [13] 87.5 114.8 91.1 157.2 84.6 116.9 63.2 265.3 81.6 156.3 +GC-TTT [38] 85.2 121.9 87.9 165.8 80.7 123.6 55.9 277.1 77.4 162.8 +EFN (Vol=300) 88.5 115.4 91.3 158.8 85.7 117.0 72.1 267.2 84.4 160.0 +EFN (Vol=1000) 89.9 109.0 92.2 156.1 89.2 115.2 76.5 261.9 87.0 156.7 Base: UniVLA UniVLA [8] 96.5 112.7 96.8 159.0 95.6 124.9 92.0 264.5 95.2 164.2 +kNN-RAG 94.2 118.3 95.7 165.1 93.1 128.6 87.5 270.3 92.6 168.9 +ResAct [29] 97.1 109.6 97.4 154.7 96.2 122.1 89.8 260.7 95.1 160.0 +R2A [13] 97.5 107.2 97.8 152.3 96.6 120.5 91.3 257.9 95.8 158.0 +GC-TTT [38] 96.8 110.1 97.1 156.8 95.9 123.4 89.1 266.1 94.7 162.1 +EFN (Vol=300) 97.7 103.4 97.9 151.4 97.2 120.1 93.7 253.8 96.6 156.2 +EFN (Vol=1000) 98.2 102.1 98.2 145.8 97.6 117.8 94.6 242.5 97.2 151.3 Base: GO-1 GO-1 [7] 96.3 111.6 97.4 160.2 95.6 118.9 89.3 269.1 94.7 163.1 +kNN-RAG 94.1 116.8 95.9 167.5 93.4 122.7 85.1 274.9 92.1 168.0 +ResAct [29] 96.8 108.9 97.6 156.1 96.1 119.8 88.7 264.7 94.8 160.3 +R2A [13] 97.3 106.5 97.9 153.6 96.5 118.2 90.2 261.5 95.5 158.1 +GC-TTT [38] 96.6 113.2 97.2 159.5 95.8 121.1 88.9 267.3 94.6 163.3 +EFN (Vol=300) 97.5 107.8 98.1 154.5 96.8 115.6 91.2 258.5 95.9 157.5 +EFN (Vol=1000) 98.1 105.3 98.5 151.7 97.3 114.1 92.8 253.2 96.7 154.8
@@
Figure 6: Visualization of EFN’s residual corrections. Left: current observation. Middle: retrieved frame. Right: corrected (translucent) versus base action. Even when the retrieved scene differs from the current one (top row), EFN can still extract actionable guidance.
@@
Table 2: Real-world performance on the AgiBot-G1 platform with the GO-1 policy [7].
@@
Method BottlePlace ShelfSort StockLift DrawerStore Average Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ GO-1 [7] 47.3 407.0 34.0 442.1 16.0 389.2 5.3 728.9 25.8 491.8 +kNN-RAG 35.3 435.2 25.3 468.7 12.0 405.1 0.0 752.3 18.2 515.3 +ResAct [29] 52.7 418.9 41.3 431.4 22.7 396.8 8.0 715.6 31.2 490.7 +R2A [13] 61.3 399.1 51.3 415.7 31.3 372.5 14.0 689.2 39.5 469.1 +GC-TTT [38] 49.3 412.5 36.7 448.9 18.7 393.7 0.0 735.1 26.2 497.6 +EFN (Vol=300) 69.3 382.6 54.7 410.5 42.0 355.1 37.3 668.3 50.8 454.1 +EFN (Vol=1000) 82.0 364.2 74.7 392.9 65.3 339.9 58.7 643.4 70.2 435.1
@@ 4 Experiments
We evaluate EFN on the LIBERO benchmark [25] with three pretrained VLAs: OpenVLA [18], UniVLA [8], and GO-1 [7]. All models use the same visual preprocessing (center crop followed by resize to $256$). UniVLA exposes $256$ visual tokens and $4$ latent action tokens per step; EFN takes the current tokens and retrieved tokens, predicts a residual in latent space, and adds it to the base latent action, which is then decoded by the frozen action head. Training details are provided in Appendix SectionB.
Evaluation protocol. At test time, we make the policy deterministic by applying $\tanh$ to the actor mean. We report success rate and, conditional on success, the average number of steps (lower is better) with horizon H=320H{=}320. The experience bank used at evaluation mirrors training and is queried via cosine nearest neighbors over pooled visual embeddings. After each successful rollout, we append the episode to the bank and prioritize shorter successful episodes; no gradient updates are performed during evaluation.
Baselines and variants. We compare each backbone with four baselines and our EFN variants, chosen to instantiate the three paradigms from Sec. 2. kNN-RAG is a non-parametric retrieval-only controller that retrieves the nearest past step from the shared bank and directly copies its action. ResAct [29] is a residual-only baseline that learns an additive correction on top of the frozen VLA action but does not use retrieval or similarity-shaped rewards. R2A [13] adapts retrieval-augmented RL to VLAs: the policy is conditioned on retrieved trajectories and optimized end-to-end, updating the backbone. GC-TTT [38] is a goal-conditioned test-time training baseline that updates the backbone online using retrieved rollouts. Additional baseline details are provided in Appendix SectionB.4.
@@ 4.2 Benchmark and Analysis
Real-world results. Table 2 reports results on AgiBot-G1 with GO-1 on four real-world tasks: BottlePlace, ShelfSort, StockLift, and DrawerStore, ordered roughly from easy to hard. Detailed settings for these tasks are provided in Appendix SectionB.5. The trends mirror simulation: kNN-RAG consistently hurts performance; ResAct and R2A offer moderate gains; and GC-TTT is unstable and fails on the hardest task. EFN achieves the highest average success with fewer steps, and the gap between Vol=300 and Vol=1000 is smaller than in simulation, suggesting that a moderate bank already covers the main task variations. All EFN results are obtained with a frozen GO-1 backbone; adaptation arises solely from the experience bank and the residual head, whereas R2A and GC-TTT rely on gradient-based updates. We include visualizations of the real-world experiments in Appendix SectionF.
@@ 4.3 Ablation Study
We further evaluate four controlled variants on LIBERO (Figure 7); full ablation numbers are provided in Appendix Table D.1.
Effect of SAC. Removing SAC (w/o SAC) and replacing it with a value-only critic harms both success and efficiency, supporting the use of entropy-regularized optimization.
@@
Figure 7: Ablation studies of EFN on the LIBERO benchmark.
Table 3: Inference-time efficiency comparison for real-world deployment. Change is computed relative to the base GO-1.
@@
Effect of instruction. Disabling instruction-level filtering (w/o instr) degrades performance, indicating that instruction-aware retrieval is important.
Effect of anti-idle. Removing the anti-idling term (w/o anti-idle) reduces performance, indicating that idleness penalties matter.
Effect of residual action. As further analyzed in Appendix SectionE.2, replacing the residual correction with direct action prediction degrades performance, which we attribute to discarding the base policy’s generalization and overburdening the lightweight actor with modeling the full action space.
Inference time. EFN improves performance while remaining efficient. Although each step introduces a modest $4.2\%$ latency overhead, EFN reduces the average episode length by $11.5\%$, resulting in an overall $7.9\%$ reduction in total episode time. In other words, EFN achieves higher success rates while also reducing end-to-end episode time.
@@ 5 Conclusion
Inspired by the human experience of déjà vu, we introduce the Experience Feedback Network, a retrieval-conditioned residual module that augments a frozen vision–language–action policy with an experience bank. EFN retrieves task-relevant transitions in a joint vision–language space and trains a lightweight residual controller with similarity-shaped reinforcement learning, enabling post-deployment adaptation through memory updates instead of backbone finetuning. On LIBERO and a real-world AgiBot-G1 platform, EFN consistently outperforms retrieval-only, residual-only, retrieval-augmented RL, and test-time training baselines under the same interaction budget.
Acknowledgments
This research is supported by the Fundamental Research Funds for the Central Universities (project number YG2024ZD06), NSFC (No. 62176155), and Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102).
@@ Appendix A Overview
This appendix is organized as follows. Section B describes the detailed settings of our method, baselines, and experiments. Section C elaborates on the rationale, motivation, and intuition behind EFN. Section D discusses the method and its broader implications. Section E presents supplementary experimental results and analysis. Section F provides additional visualizations of EFN in action.
@@ Appendix B Details of Method and Experiment
This section provides implementation and experimental details; the rationale behind these design choices is discussed in Section C.
@@ B.1 Experience Bank Implementation Details
$E_{m,n}=\big(\ell_{\tau_{m}},\,\mathbf{F}_{m,n},\,\mathbf{k}_{m,n},\,\mathbf{a}^{(0)}_{m,n}\big)$, (B.1)
where $\ell_{\tau_{m}}$ is the instruction embedding for rollout $\tau_{m}$, $\mathbf{F}_{m,n}\in\mathbb{R}^{L\times C}$ is the VLA vision encoder feature map for frame $s_{m,n}$, $\mathbf{k}_{m,n}\in\mathbb{R}^{d_{k}}$ is the mean–max fused retrieval key defined in main-paper Sec. 3.2, and $\mathbf{a}^{(0)}_{m,n}$ is the raw action produced by the frozen base policy at that step.
In our implementation, the bank is stored as a line-based JSON file together with accompanying feature files, where each JSON line encodes one step and includes a rollout identifier, a step index, and paths to the visual and action features. The corresponding visual feature file stores the precomputed retrieval key $\mathbf{k}_{m,n}$ as a $d_{k}$-dimensional vector (e.g., visual_embed_meanmax with $d_{k}=4096$), together with any additional visual features that may be needed. The action file stores the base policy’s latent action representation (e.g., a $4\times d_{a}$ tensor of action tokens) and the associated discrete token IDs (generated_ids) from which the continuous control command $\mathbf{a}^{(0)}_{m,n}$ is decoded.
Figure B.1: Cosine similarities (left bars) and corresponding softmax sampling probabilities (right bars) over the top candidates in the experience bank. Higher similarity induces higher sampling probability while still preserving exploration.
@@
$\texttt{index}:(\texttt{rollout\_id},\,\texttt{step\_idx})\mapsto E_{m,n},$ (B.2)
@@
$\mathbf{K}\in\mathbb{R}^{N\times d_{k}},$ (B.3)
@@
$E_{m,n+1}=\texttt{index}\big(\texttt{rollout\_id}=m,\;\texttt{step\_idx}=n+1\big).$ (B.4)
@@
Given a current observation $s_{t}$, we compute the vision encoder feature map $\mathbf{F}_{t}$ and derive a query key $\mathbf{q}_{t}$ via the same mean–max fusion and $\ell_{2}$-normalization used for $\mathbf{k}_{m,n}$ (main-paper Sec. 3.2). We then compute cosine similarities between $\mathbf{q}_{t}$ and all stored keys:
$s_{i}=\cos(\mathbf{q}_{t},\mathbf{k}_{i}),\quad i\in\{1,\dots,N\}.$ (B.5)
In the probabilistic variant used for EFN training, we follow the top-$k$ sampling scheme described in the main text: we first select a shortlist $\mathcal{N}_{k}(\mathbf{q}_{t})$ of the $k$ highest-similarity keys and sample an index using a temperature-scaled softmax,
$p(i\mid\mathbf{q}_{t})=\frac{\exp\big(s_{i}/\tau\big)}{\sum_{j\in\mathcal{N}_{k}(\mathbf{q}_{t})}\exp\big(s_{j}/\tau\big)},\quad i\in\mathcal{N}_{k}(\mathbf{q}_{t}),$ (B.6)
with a fixed temperature $\tau>0$ (we use $\tau=0.05$ in all experiments). As shown in Figure B.1, this “retrieve-then-sample” design encourages exploration among near-matches while still biasing toward the most semantically similar experiences.
@@
During offline pretraining of EFN, the experience bank is constructed from a mixture of successful, near-successful, and occasionally failed rollouts, as discussed in Appendix Section D. At each environment step, EFN queries the bank, obtains a reference step $E_{m,n}$ and its successor $E_{m,n+1}$, and uses the successor observation as the semantic target for the similarity-shaped progress reward. At deployment, the bank continues to grow: we append only the trajectories that reach the goal, thereby gradually refining the set of available references without ever updating the frozen VLA backbone weights.